OpenStudy (anonymous):

The line that passes through the given point and is perpendicular to the given line in slope intercept form: (1,-9) 4x+8y=12

OpenStudy (anonymous):

First, rewrite the line in slope intercept form.

OpenStudy (anonymous):

Okay will it be um y=2x+3/2 ??? im not positive about it

OpenStudy (anonymous):

Almost. Divide everything in the equation by 4 first.

OpenStudy (anonymous):

And then isolate y.

OpenStudy (anonymous):

Ok I did I got y=1/2x+3/2

OpenStudy (anonymous):

Yup. So one trick with perpendicular equations is that the slope is always the negative reciprocal of the current slope.

OpenStudy (anonymous):

So if your slope is 2, then the perpendicular slope is -1/2.

OpenStudy (anonymous):

So in this case, your slope is 1/2, which means the perpendicular slope is -2.

OpenStudy (anonymous):

Got it and is the slope always going to come up as a fraction?

OpenStudy (anonymous):

Not necessarily, it depends on what your original slope was. If your original slope was a fraction, then the perpendicular slope may be an integer or a fraction, depending on what the fraction was. i.e. 10/7 would become -7/10 but 1/10 would become -10 And we know that the slope intercept form of a line is y = mx +b. Now we have the slope as well as a point. So substitute everything into the equation to find b.

OpenStudy (anonymous):

b=-19/2

OpenStudy (anonymous):

So now you know your slope and the value of b. All you have to do is put it in y = mx + b form!

OpenStudy (anonymous):

so then the slope intercept form will be y=1/2x-19/2 ??

OpenStudy (anonymous):

am i right?

OpenStudy (anonymous):

Almost. Your original slope was 1/2, right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

So the slope of the perpendicular line will be the negative reciprocal, which means it will be -2.

OpenStudy (anonymous):

that'll mean it will y=-2-19/2 ??

OpenStudy (anonymous):

You forgot the x, but yes.

OpenStudy (anonymous):

WHOOHOO!! haha:) and lets say the slope will be -2 then what will the reciprocal be?

OpenStudy (anonymous):

1/2

OpenStudy (anonymous):

oh the reciprocal? The reciprocal is just -1/2

OpenStudy (anonymous):

if you have a value n, then the reciprocal is just 1/n

OpenStudy (anonymous):

thanks will you mind helping me in another problem?

OpenStudy (anonymous):

I'll write it in another post though